OpenStudy (anonymous):

Prove - tan^2(x) + sec^2(x) = 1 by working on one side to match the other using identities.

OpenStudy (anonymous):

\[- \tan^2x + \sec^2x = 1\]

OpenStudy (anonymous):

\[- \tan^2x + \sec^2x = 1\]

Nnesha (nnesha):

sec^2 theta = what ? remember the identity ?

OpenStudy (anonymous):

1/cos^2x

Nnesha (nnesha):

well that's reciprocal of sec but it's okay we can use that too!! tan^2 =what ?

OpenStudy (anonymous):

sin^2x/cos^2x

OpenStudy (anonymous):

or 1/cot^2x

Nnesha (nnesha):

yes right so replace tan and sec with that \[\huge\rm -\frac{ \sin^2x }{ \cos^2x } +\frac{ 1 }{ \cos^2x}\] find the common denominator

OpenStudy (anonymous):

cos^2x?

Nnesha (nnesha):

ohh well not gonna work should use the identity i guess

OpenStudy (jhannybean):

You could also use the fact that \(\sf sec^2(\theta) = tan^2(\theta) +1\) and then substitute this in place of \(\sf \sec^2(\theta)\)

OpenStudy (anonymous):

thats true, thanks

OpenStudy (jhannybean):

\[\sf -tan^2(\theta)+\sec^2(\theta) = 1\]\[\sf -\tan^2(\theta) +\color{red}{\tan^2(\theta) +1}=1\]

OpenStudy (anonymous):

Wow... The one identity I didn't think of solved it so easily. Thank you!

Nnesha (nnesha):

\[\huge\rm \frac{ -\sin^2x +1}{ \cos^2 }\] use the special identity sin^2x+cos^2x =1 solve for cos^2

OpenStudy (jhannybean):

No problem :)

Nnesha (nnesha):

here you can copy these identities http://www.math.com/tables/trig/identities.htm you weren't familiar with this so that's why i thought better to write interms of sin and cos

OpenStudy (jhannybean):

That's a good way to approach it too. @Nnesha :)

OpenStudy (anonymous):

Thanks that will help too @Nnesha

Nnesha (nnesha):

yw :=)